OpenStudy (anonymous):

Quick help on Algebra 1 "The line of best fit" Write the equation of the line passing through those two points using the point-slope formula y - y1 = m(x - x1). Show all of your work. Remember to find the slope of the line first. The two points are: (43.2,13),(67.2,21)

OpenStudy (anonymous):

start by finding the slope \[\frac{21-13}{67.2-43.2}\]

OpenStudy (anonymous):

Okay one second

OpenStudy (anonymous):

8/24

OpenStudy (anonymous):

i get \(\frac{8}{24}=\frac{1}{3}\) i think that is right

OpenStudy (anonymous):

then go to \[y-21=\frac{1}{3}(x-67.2)\]

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

unless you want to multiply out and solve for \(y\)

OpenStudy (anonymous):

Nope! Thanks :)

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

depends on the form you want to put the line it it is not that hard, use a calculator

OpenStudy (anonymous):

@satellite73 i just reposted the question from before

OpenStudy (anonymous):

How'd you go from 8/24=1/2 to y-21...

OpenStudy (anonymous):

there is no \(\frac{1}{2}\) it is \(\frac{8}{24}=\frac{1}{3}\) by reducing the fraction that is the slope \(m\) then use the point slope formula that you wrote above

OpenStudy (anonymous):

that is why you get \[y-21=\frac{1}{3}(x-67.2)\] \[y-21=\frac{1}{3}x-22.4\] \[y=\frac{1}{3}x-1.4\]

OpenStudy (anonymous):

and that ^ is the final answer?

OpenStudy (anonymous):

that should do it, yes

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

yw